OpenStudy (anonymous):

How many hydrogen atoms are contained in 1.00 g of hydrogen atoms? In 1.00 g of hydrogen molecules?

OpenStudy (anonymous):

Step 1: Convert mass of hydrogen atoms to moles. Divide by the atomic weight. Step 2: Convert the number of moles obtained in Step 1) to particles by multiplying by Avogadro's number. For the second part of your question, hydrogen gas is diatomic, so multiply the answer you get above by two.

OpenStudy (thadyoung):

You will need avogadro's number. First, you need to divide by the molar mass of 1 hydrogen atom, which you can find on the periodic table. If you don't have one, you can use this one: http://www.ptable.com/ Then, multiply by avogadro's number, which is some unit per mole, you will see that this is simply algebra, you're canceling out units: avogadro's number is: \(\sf \color{blue}{6.022 \times 10^{23}} \color{red}{\frac{units}{mole}} \)

OpenStudy (thadyoung):

\(\sf \cancel{mole} \times \large \frac{atoms }{\cancel{mole}} = atoms \)